OpenStudy (anonymous):

Find the exact value of sin-1the quantity square root of two divided by two.

OpenStudy (anonymous):

\[\sin^{-1} (\sqrt{2/}2)\]

OpenStudy (anonymous):

@welshfella

rishavraj (rishavraj):

is it \[\sqrt{\frac{ 2 }{ 2 }}~~or~~\frac{ \sqrt{2} }{ 2 }\] if its \[\frac{ \sqrt{2} }{ 2 } ~~then ~~\it~~means~~tht~~\frac{ 1 }{ \sqrt{2} }\]

OpenStudy (anonymous):

it has to be in radians @rishavraj

rishavraj (rishavraj):

\[remember????~~~~ \sin^{-1} \frac{ \pi }{ 4 } = \frac{ 1 }{ \sqrt{2} }\]

OpenStudy (anonymous):

okay thank you

rishavraj (rishavraj):

yw :)))